# Kai Shu, Lu Cheng, Huan Liu

### v3

@@ 7.2 Dynamic & Complex Judgment
Earlier works about LLM-as-a-judge usually adopt static and straightforward approaches that directly prompt the judge LLM to perform the assessment Zheng et al. (2023). Recently, more dynamic and complex judgment pipelines have been proposed to address various limitations, improving the robustness and effectiveness of LLM-as-a-judge. One approach in direction is to follow the concept of “LLM-as-a-examiner”, where the system dynamically and interactively generates both questions and judgments based on the candidate LLMs’ performance Yu et al. (2024b); Bai et al. (2023a). Other works focus on making judgments based on two or more candidate LLMs battle and debating results Moniri et al. (2024); Zhao et al. (2024c). These dynamic judgment methods largely improve the judge LLM’s understanding of each candidate and potentially prevent the data contamination problem Deng et al. (2024); Golchin and Surdeanu in LLMs’ evaluation. Additionally, building complex and complicated judgment pipelines or agents is another popular research area Li et al. (2023); Chan et al. (2023); Zhuge et al. (2024). These methods typically involve multi-agent collaboration, along with well-designed planning and memory systems, enabling the judge LLMs to handle more complex and diverse judgment scenarios.
@@ 7.3 Self-Judging
LLM-based evaluators, such as GPT-4, are widely used for assessing outputs but face significant challenges, particularly Egocentric Bias, where models favor their own responses over those from external systems Liu et al. (2023b); Bai et al. (2023a); Zheng et al. (2023). This self-preference undermines impartiality and creates a “chicken or the egg” dilemma: robust evaluators are essential for developing powerful LLMs, yet advancing LLMs depends on unbiased evaluators. Other issues include Self-Enhancement Bias, where models overrate their own outputs Li et al. (2023), and Reward Hacking, where over-optimization of specific signals leads to less generalizable evaluations Wu et al. (2024a). Additionally, reliance on Static Reward Models limits adaptability, while biases like Positional and Verbosity Biases distort judgments by favoring response order or length over quality Yuan et al. (2024c); Wang et al. (2024f). The high cost and limited scalability of human annotations further complicate the creation of dynamic and reliable evaluation systems Srivastava et al. (2022); Liang et al. (2022).
@@
Future Direction. One promising direction for future research is the development of collaborative evaluation frameworks like Peer Rank and Discussion (PRD) Li et al. (2023). These frameworks leverage multiple LLMs to collectively evaluate outputs, using weighted pairwise judgments and multi-turn dialogues to reduce self-enhancement bias and align evaluations closer to human standards. Another intriguing avenue is the adoption of Self-Taught Evaluator frameworks, which generate synthetic preference pairs and reasoning traces to iteratively refine model evaluation capabilities Wang et al. (2024f). This approach eliminates dependency on costly human annotations while ensuring that evaluation criteria adapt to evolving tasks and models. The integration of Self-Rewarding Language Models (SRLM) offers another promising path Yuan et al. (2024c). By employing iterative mechanisms like Direct Preference Optimization (DPO), these models continuously improve their instruction-following and reward-modeling capabilities, mitigating issues like reward hacking and overfitting. Building on SRLM, the use of Meta-Rewarding mechanisms introduces a meta-judge role to evaluate and refine judgment quality Wu et al. (2024a). This iterative process addresses biases such as verbosity and positional bias, enhancing both alignment and the ability to evaluate complex tasks. Finally, leveraging synthetic data creation for generating contrasting responses offers a scalable solution for training evaluators Wang et al. (2024f). By iteratively refining evaluations on synthetic preference pairs, models can progressively improve their robustness and adaptability. Combining these approaches with diverse benchmarks Srivastava et al. (2022); Liang et al. (2022), multifaceted evaluation criteria Sellam et al. (2020), and human feedback Xu et al. (2023); Ouyang et al. (2022) can ensure evaluations that are fair, reliable, and aligned with human expectations across various domains.

### v4

@@ 7.2 Dynamic & Complex Judgment
Earlier works about LLM-as-a-judge usually adopt static and straightforward approaches that directly prompt the judge LLM to perform the assessment Zheng et al. (2023). Recently, more dynamic and complex judgment pipelines have been proposed to address various limitations, improving the robustness and effectiveness of LLM-as-a-judge. One approach in direction is to follow the concept of “LLM-as-a-examiner”, where the system dynamically and interactively generates both questions and judgments based on the candidate LLMs’ performance Yu et al. (2024c); Bai et al. (2023a); Kim et al. (2024a). Other works focus on making judgments based on two or more candidate LLMs battle and debating results Moniri et al. (2024); Zhao et al. (2024c); Kenton et al. . These dynamic judgment methods largely improve the judge LLM’s understanding of each candidate and potentially prevent the data contamination problem Deng et al. (2024); Golchin and Surdeanu in LLMs’ evaluation. Additionally, building complex and complicated judgment pipelines or agents is another popular research area Li et al. (2023a); Chan et al. (2023); Zhuge et al. (2024); Zheng et al. (2024). These methods typically involve multi-agent collaboration, along with well-designed planning and memory systems, enabling the judge LLMs to handle more complex and diverse judgment scenarios.
@@ 7.3 Self-Judging
LLM-based evaluators, such as GPT-4, are widely used for assessing outputs but face significant challenges, particularly Egocentric Bias, where models favor their own responses over those from external systems Liu et al. (2023b); Bai et al. (2023a); Zheng et al. (2023). This self-preference undermines impartiality and creates a “chicken or the egg” dilemma: robust evaluators are essential for developing powerful LLMs, yet advancing LLMs depends on unbiased evaluators. Other issues include Self-Enhancement Bias, where models overrate their own outputs Li et al. (2023a), and Reward Hacking, where over-optimization of specific signals leads to less generalizable evaluations Wu et al. (2024a). Additionally, reliance on Static Reward Models limits adaptability, while biases like Positional and Verbosity Biases distort judgments by favoring response order or length over quality Yuan et al. (2024c); Wang et al. (2024f). The high cost and limited scalability of human annotations further complicate the creation of dynamic and reliable evaluation systems Srivastava et al. (2022); Liang et al. (2022).
@@
Future Direction. One promising direction for future research is the development of collaborative evaluation frameworks like Peer Rank and Discussion (PRD) Li et al. (2023a). These frameworks leverage multiple LLMs to collectively evaluate outputs, using weighted pairwise judgments and multi-turn dialogues to reduce self-enhancement bias and align evaluations closer to human standards. Another intriguing avenue is the adoption of Self-Taught Evaluator frameworks, which generate synthetic preference pairs and reasoning traces to iteratively refine model evaluation capabilities Wang et al. (2024f). This approach eliminates dependency on costly human annotations while ensuring that evaluation criteria adapt to evolving tasks and models. The integration of Self-Rewarding Language Models (SRLM) offers another promising path Yuan et al. (2024c). By employing iterative mechanisms like Direct Preference Optimization (DPO), these models continuously improve their instruction-following and reward-modeling capabilities, mitigating issues like reward hacking and overfitting. Building on SRLM, the use of Meta-Rewarding mechanisms introduces a meta-judge role to evaluate and refine judgment quality Wu et al. (2024a). This iterative process addresses biases such as verbosity and positional bias, enhancing both alignment and the ability to evaluate complex tasks. Finally, leveraging synthetic data creation for generating contrasting responses offers a scalable solution for training evaluators Wang et al. (2024f). By iteratively refining evaluations on synthetic preference pairs, models can progressively improve their robustness and adaptability. Combining these approaches with diverse benchmarks Srivastava et al. (2022); Liang et al. (2022), multifaceted evaluation criteria Sellam et al. (2020), and human feedback Xu et al. (2023); Ouyang et al. (2022) can ensure evaluations that are fair, reliable, and aligned with human expectations across various domains.